# SplatFormer: Point Transformer for Robust 3D Gaussian Splatting

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has recently transformed photorealistic reconstruction, achieving high visual fidelity and real-time performance. However, rendering quality significantly deteriorates when test views deviate from the camera angles used during training, posing a major challenge for applications in immersive free-viewpoint rendering and navigation. In this work, we conduct a comprehensive evaluation of 3DGS and related novel view synthesis methods under out-of-distribution (OOD) test camera scenarios. By creating diverse test cases with synthetic and real-world datasets, we demonstrate that most existing methods, including those incorporating various regularization techniques and data-driven priors, struggle to generalize effectively to OOD views. To address this limitation, we introduce SplatFormer, the first point transformer model specifically designed to operate on Gaussian splats. SplatFormer takes as input an initial 3DGS set optimized under limited training views and refines it in a single forward pass, effectively removing potential artifacts in OOD test views. To our knowledge, this is the first successful application of point transformers directly on 3DGS sets, surpassing the limitations of previous multi-scene training methods, which could handle only a restricted number of input views during inference. Our model significantly improves rendering quality under extreme novel views, achieving state-of-the-art performance in these challenging scenarios and outperforming various 3DGS regularization techniques, multi-scene models tailored for sparse view synthesis, and diffusion-based frameworks. Project page: https://sergeyprokudin.github.io/splatformer/.
@@ 1 Introduction
Novel view synthesis (NVS) focuses on transforming 2D RGB images into immersive 3D scenes, allowing users to navigate within augmented reality (AR) and virtual reality (VR) environments. Traditionally, this problem has been approached using a standard novel view interpolation protocol, where test views are sampled at fixed intervals along the trajectory of the input views. Several NVS methods have emerged based on this protocol, with 3D Gaussian splatting (3DGS) (Kerbl et al., 2023) recently gaining attention for achieving real-time and high-fidelity results in view interpolation.
However, AR and VR applications require not only smooth transitions between input views but also the ability to explore novel regions of interest from viewing angles outside the input distribution. For instance, users may want to observe a scene from high-elevation angles, which are often missing from the captured views. While novel view interpolation has seen significant advancements, this out-of-distribution novel view synthesis (OOD-NVS) task remains underexplored, particularly in terms of evaluation protocols and methodology. A related area of research involves 3D reconstruction from sparse or monocular input views, where methods often need to hallucinate unseen content (Liu et al., 2023a; Chan et al., 2023; Kwak et al., 2024; Liu et al., 2023b). While hallucination can be beneficial for creative applications, it may be undesirable in settings that demand accurate reconstructions, such as 3D visualization of surgical procedures (Hein et al., 2024), and unnecessary in typical daily capture scenarios.
@@
Figure 1: We investigate the out-of-distribution (OOD) novel view synthesis (NVS), where test views significantly differ from input views. This scenario contrasts with prior in-distribution NVS, where test views interpolate between densely captured input views, Sparse NVS with a few large-baseline input views, and Nerfbusters NVS (Warburg* et al., 2023), where test views share similar angles with input views. Existing NVS methods, including MipNeRF360 (Barron et al., 2022), and those designed for sparse inputs like LaRa (Chen et al., 2024a), face challenges in this setting, while our method shows notable improvements.
@@
Unlike previous works that rely on hand-crafted regularization techniques (Xie et al., 2024; Li et al., 2024b), we adapt point transformer (Zhao et al., 2021), an attention-based architecture designed for 3D scene understanding, to process the 3DGS as a point cloud set with Gaussian attributes serving as features. The attention mechanism in the point transformer learns to capture multi-view information embedded in the 3DGS, focusing on the local neighborhood within the spatial structure pre-computed by the initial 3DGS. It outputs residuals that are added to the input Gaussian attributes. The updated 3DGS is then rendered from novel views, and a photometric error between the rendered and ground-truth images is minimized to train the SplatFormer. We curate large-scale training pairs of initial, flawed 3DGS sets, and ground-truth images of in-distribution and OOD views using ShapeNet and Objaverse 1.0, which are made feasible by the fast optimization of 3DGS and the availability of large-scale 3D and multi-view datasets. By training on this dataset, SplatFormer learns generalizable priors for refining 3DGS, effectively removing artifacts in the OOD views while maintaining 3D consistency.
@@ 2 Related Work
In this section, we review several novel view synthesis approaches relevant to our topic. These methods can be broadly categorized into three modalities: (1) general NVS methods primarily evaluated through view interpolation, (2) reconstruction from sparse views, and (3) approaches that enhance NVS by utilizing priors and regularization techniques.
@@
Novel View Interpolation. In most NVS scenarios, both input and test views are sampled from the same distribution, typically following a fixed trajectory or a hemispherical pattern, as observed in common datasets such as Blender NeRF (Mildenhall et al., 2020), LLFF (Mildenhall et al., 2019), and Phototourism (Jin et al., 2021). Seminal works like NeRF (Mildenhall et al., 2020), InstantNGP (Müller et al., 2022), and 3DGS (Kerbl et al., 2023) have demonstrated strong performance under this view interpolation protocol. However, as we will demonstrate later, they encounter difficulties in rendering novel views from out-of-distribution (OOD) angles, a challenge that remains less explored.
@@
Sparse View Reconstruction. Another line of research is focused on reconstructing and rendering 3D scenes from sparse set of input views, typically fewer than four. Methods like Zero123 (Liu et al., 2023a) and others (Sargent et al., 2024; Liu et al., 2023b; Kwak et al., 2024) predict NeRF representations or multi-view images from monocular inputs, often requiring the model to hallucinate unseen regions. Similarly, approaches like (Yu et al., 2021; Charatan et al., 2024; Wewer et al., 2024; Chen et al., 2024a; b) predict 3D primitives directly from sparse views. Some methods (Höllein et al., 2024; Gao et al., 2024) aggregate multi-view inputs into 2D diffusion-based generative models, which are more suited for creative tasks but less applicable when metrically accurate reconstructions are needed.
@@
Out-of-distribution Novel View Synthesis (OOD-NVS). Unlike the standard interpolation setup, Nerfbusters (Warburg* et al., 2023) captures input and test views along separate trajectories, closely aligning with the OOD-NVS problem we are addressing. However, their input and test views remain relatively similar in viewing angles, and the observed artifacts are primarily caused by the “invisibility issue“, where test views fall outside the input observation sphere, rather than from significant viewpoint deviations. In contrast, our approach tackles large viewpoint shifts without addressing invisibility, filling a gap by emphasizing generalization across substantial angle deviations.
@@
Regularization Techniques for Unconstrained Reconstruction. Sparse input views significantly degrade NVS performance, leading to various works exploring geometric priors, spatial regularity constraints, and data-driven priors. Geometric priors have been used in SuGaR (Guédon & Lepetit, 2024), 2DGS (Huang et al., 2024a), and GeoGaussian (Li et al., 2024b), which apply handcrafted self-supervision losses to better align Gaussian splats with surface geometry. Similarly, Spatial regularity constraints, explored in SplatFields (Mihajlovic et al., 2024) and ZeroRF (Shi et al., 2024), integrate Deep Image Prior (Ulyanov et al., 2018) to regularize 3DGS and NeRF reconstructions, producing more robust results from sparse inputs. However, these methods offer limited improvements as they do not leverage external data. Data-driven priors have been adopted in several works. FSGS (Zhu et al., 2024) and DNGaussian (Li et al., 2024a) supervises depth maps using deep stereo models but suffers from scale ambiguity. InstantSplat (Fan et al., 2024) uses dense point clouds for 3DGS initialization, though it struggles with overfitting. Nerfbusters (Warburg* et al., 2023) pretrains a diffusion model for post-processing NeRF, achieving only marginal improvements. Appearance priors (Zhu et al., 2024; Sargent et al., 2024; Wu et al., 2024a; Gao et al., 2024; Kwak et al., 2024) use 2D diffusion models to regularize novel view renderings, but often struggle with multi-view consistency. Additionally, SSDNeRF (Chen et al., 2023) and HypNeRF (Sen et al., 2023) pretrain 3D feature grids on object-centric datasets, yet underperform in multi-category experiments.
@@
Learning-based 2D-to-3D Models. Another special case of data-driven models involves training feed-forward models on large-scale multi-view image datasets to predict 3D representations from 2D images. SyncDreamer (Liu et al., 2023b) and ViewDiff (Höllein et al., 2024) fine-tune pretrained diffusion models to generate multi-view images from one or a few input views. PixelNeRF (Yu et al., 2021), MVSplat (Chen et al., 2024b), and related works (Charatan et al., 2024; Wewer et al., 2024) transform 2D image features into NeRF or Gaussian splats. Although these models can learn generic priors from multi-view datasets, they are typically constrained to only a few input views, limiting their ability to fully leverage larger multi-view inputs.
@@
3D Point Processing Techniques are central to our work and widely used across 3D tasks. Unlike 2D image features or 3D grids, point clouds are unordered and unevenly distributed, requiring specialized architectures to handle their irregularity and sparsity. Solutions include sparse convolution (Choy et al., 2019), MLPs (Qi et al., 2017), and transformers (Wang, 2023; Yang et al., 2023; Zhao et al., 2021). The point transformer (Zhao et al., 2021), using attention to model spatial relationships, has proven particularly effective. Further advancements (Wu et al., 2022; 2024b) have enhanced its efficiency, achieving state-of-the-art performance in 3D segmentation. Our work is the first to adapt the point transformer for refining 3D Gaussian splats (3DGS), leveraging its ability to capture spatial relationships in irregular point clouds to enhance novel view synthesis fidelity.
@@ 4 Robust Out-of-distribution Novel View Synthesis
Key Observation. As shown in Fig. 2, the reconstruction quality degrades significantly as the test camera’s elevation increases, highlighting a key limitation of 3DGS in handling OOD views. The challenge is to make the representation robust to such viewpoint changes while preserving the advantages of 3DGS, such as real-time rendering and compatibility with rasterization-based tools. Addressing this limitation by incorporating priors and constraints into the optimization of 3DGS is a complex task. Previous approaches have attempted to address this using geometric constraints (Huang et al., 2024a; Mihajlovic et al., 2024) and data-driven priors (Fan et al., 2024). However, as demonstrated later (Tab. 1), these methods fall short in achieving robust novel-view synthesis, emphasizing the need for a more effective solution. We believe that solving this issue requires incorporating three key aspects: leveraging generic priors from large-scale datasets, ensuring 3D consistency in renderings, and fully utilizing rich geometric information from all input views.
@@ 4.1 Learning Data-Driven Prior
Dataset. To enable SplatFormer to refine imperfect Gaussian splats using a data-driven prior, we curated a large dataset containing pairs of Gaussian primitives and corresponding multi-view images. We utilized 33k and 48k scenes from the ShapeNet (Chang et al., 2015) and Objaverse-1.0 (Deitke et al., 2023) datasets respectively. These assets were rendered from low-elevation input views and high-elevation out-of-distribution (OOD) views. The initial splats were generated from the low-elevation views (following Sec. 3). The data collection process, which required approximately 3000 GPU hours, was efficiently executed using budget GPUs like the RTX-2080Ti. We will release the data and corresponding rendering code to facilitate future research.
@@
Figure 4: Novel View Synthesis under Out-of-Distribution Camera Angles. The first column shows 4 out of 32 input views. Here, we compare our method with LaRa (Chen et al., 2024a), SplatFields (Mihajlovic et al., 2024), MipNeRF360 (Barron et al., 2022), 2DGS (Huang et al., 2024a), and 3DGS (Kerbl et al., 2023). Results on Objaverse-OOD evaluation scenes; a comprehensive comparison with all the baselines is provided in the appendix (Fig. F.3).
@@ 5 Experiments
We begin by outlining the out-of-distribution novel-view synthesis (OOD-NVS) setup, followed by a description of the evaluation protocol and the baseline methods used for comparison. Next, we present the results on OOD-NVS, cross-dataset generalization, and ablation studies. Finally, we discuss the limitations of our approach and potential directions for future research.
@@
OOD-NVS. For a given region of interest, such as a centered object, we simulate an input camera capturing 360-degree azimuths at low elevations. The camera takes $N_{\textrm{in}}$ photos from evenly spaced azimuths, with its elevation following a sinusoidal pattern defined by frequency $f$ and limits $\phi_{\textrm{max}}$. This setup mimics a user recording the target from low angles, with physical constraints preventing stable top-down captures. To introduce OOD views, we use high-elevation with $\phi_{\textrm{ood}}\gg\phi_{\textrm{max}}$, simulating a top-down perspective. The camera configuration is visualized in Fig 1.
@@
Evaluation Sets. We use Blender to render 20 objects from ShapeNet (Chang et al., 2015) and Objaverse-v1 (Deitke et al., 2023) each. We select common objects and scenes with meaningful bird’s-eye views, such as city streets and buildings, avoiding those with large cavities that are invisible from low-elevation views. Due to varying object heights and shapes, we render two input camera trajectories with $\phi_{\textrm{max}}=10^{\circ}$ and $20^{\circ}$, creating two input-target splits to represent different levels of view deviation. So each scene has two experiments with different input views and the same target views. We average evaluation metrics across the two sets of experiments. Each input trajectory consists of $N_{\textrm{in}}=32$ views. The OOD test set includes $N_{\textrm{out}}=9$ views, uniformly distributed from the top sphere with $\phi_{\textrm{ood}}\geq 70^{\circ}$. All renderings are at a resolution of $256\times 256$.
@@
Baselines. We evaluate SplatFormer using the OOD-NVS protocol, comparing it with several publicly available state-of-the-art methods from each NVS method category. First, we include per-scene NVS approaches designed primarily for in-distribution NVS, including representative works such as InstantNGP (Müller et al., 2022), 3DGS (Kerbl et al., 2023), and MipNeRF360 (Barron et al., 2022). Next, we examine regularized 3DGS variants that aim to reduce overfitting without external priors, including 2DGS (Huang et al., 2024a) and SplatFields (Mihajlovic et al., 2024). We further consider prior-augmented per-scene NVS methods, which regularize per-scene optimization using external priors, including InstantSplat (Fan et al., 2024), FSGS (Zhu et al., 2024), SSDNeRF (Chen et al., 2023), and Nerfbusters (Warburg* et al., 2023). Finally, we examine feed-forward models, which reconstruct 2D Gaussians (LaRa (Chen et al., 2024a)) or produce multi-view images (SyncDreamer  (Liu et al., 2023b)) from sparse input images.
@@
For a fair comparison, we retrained SSDNeRF, LaRa, and SyncDreamer using the same training sets as SplatFormer. For SyncDreamer, which only supports single-view input, we selected the highest-elevation view. For LaRa, which is limited to four input views due to memory constraints, we chose four large-baseline views to maximize scene coverage; details are provided in appendix (Sec. D).
@@
Results on OOD-NVS. Qualitative results (Fig. 4) show that LaRa produces blurry outputs, while MipNeRF360 suffers from floater artifacts and SplatFields smooths out fine details. Both 2DGS and 3DGS exhibit spiky artifacts. In contrast, SplatFormer significantly reduces the artifacts present in 3DGS, completes surface reconstruction, and even restores certain geometric properties, such as interlaced structures. While our method still faces challenges with high-frequency texture details, it outperforms previous approaches in terms of fidelity and consistency in out-of-distribution views, which is also supported by the clear quantitative improvements demonstrated in our results (Tab. 1). Additional visual results are provided in the appendix (Sec. E) and the supplementary video.
@@
Figure 5: Cross-dataset Generalization. SplatFormer trained on Objaverse successfully mitigates artifacts in OOD views in the GSO (Downs et al., 2022) dataset and our real-world object-centric captures. Additional results are presented in the appendix (Fig. F.4 and Fig. F.5).
@@
On the GSO-OOD evaluation set, SplatFormer achieves substantial improvements in both metrics and visual quality. Even on the real-world dataset, despite being trained exclusively on synthetic data, SplatFormer reduces artifacts. These improvements are reflected in the SSIM and LPIPS metrics, though we observed rather minimal improvements in PSNR, which we attribute to the pixel-wise PSNR’s sensitivity to imperfect calibration and our method’s limitation in modeling specular effects. Our method also outperforms MipNeRF360 and 2DGS, the best-performing baselines in Objaverse-OOD (Tab. 1). Additionally, we evaluate Nerfbusters (Warburg* et al., 2023), which addresses robust novel view synthesis from novel camera trajectories—a relevant challenge to OOD-NVS. However, we find that Nerfbusters tends to mistakenly remove key scene content as floaters, leading to incomplete geometry.
@@
DiffBIR-stage1 (Lin et al., 2024) 24.81 0.892 0.163
DiffBIR-stage2 (Lin et al., 2024) 24.24 0.858 0.174
@@
3D vs 2D Denoising. An alternative strategy for refining OOD-NVS renderings is to use 2D image restoration methods. To explore this, we use DiffBIR (Lin et al., 2024), a state-of-the-art image restoration method, to denoise 3DGS renderings. DiffBIR consists of two cascaded models: a first-stage image-to-image regressor to remove artifacts and a second-stage diffusion-based generator (Rombach et al., 2021) to in-paint missing details. We trained both stages using pairs of OOD 3DGS renderings and ground-truth images from our ShapeNet-OOD training set. To address multi-view inconsistencies in the denoised images, we also used the generated images to retrain the 3DGS. This experiment is similar to Sp2360 (Paul et al., 2024), which uses cascaded 2D diffusion priors to regularize 3DGS from sparse-view inputs. As shown in Tab. 5, while 2D denoising methods improve the original 3DGS, they significantly underperform compared to SplatFormer and fail to recover geometric details. See the appendix (Fig. C.1) for visual comparisons.

### v3

@@ Abstract
3D Gaussian Splatting (3DGS) has recently transformed photorealistic reconstruction, achieving high visual fidelity and real-time performance. However, rendering quality significantly deteriorates when test views deviate from the camera angles used during training, posing a major challenge for applications in immersive free-viewpoint rendering and navigation. In this work, we conduct a comprehensive evaluation of 3DGS and related novel view synthesis methods under out-of-distribution (OOD) test camera scenarios. By creating diverse test cases with synthetic and real-world datasets, we demonstrate that most existing methods, including those incorporating various regularization techniques and data-driven priors, struggle to generalize effectively to OOD views. To address this limitation, we introduce SplatFormer, the first point transformer model specifically designed to operate on Gaussian splats. SplatFormer takes as input an initial 3DGS set optimized under limited training views and refines it in a single forward pass, effectively removing potential artifacts in OOD test views. To our knowledge, this is the first successful application of point transformers directly on 3DGS sets, surpassing the limitations of previous multi-scene training methods, which could handle only a restricted number of input views during inference. Our model significantly improves rendering quality under extreme novel views, achieving state-of-the-art performance in these challenging scenarios and outperforming various 3DGS regularization techniques, multi-scene models tailored for sparse view synthesis, and diffusion-based frameworks.
@@ 1 Introduction
Novel view synthesis (NVS) focuses on transforming 2D RGB images into immersive 3D scenes, allowing users to navigate in augmented reality (AR) and virtual reality (VR) environments. Traditionally, this problem has been approached using a standard novel view interpolation protocol, where test views are sampled at fixed intervals along the trajectory of the input views. Several NVS methods have emerged based on this protocol, with 3D Gaussian splatting (3DGS) (Kerbl et al., 2023) recently gaining attention for achieving real-time and high-fidelity results in view interpolation.
However, AR and VR applications require not only smooth transitions between input views but also the ability to explore novel regions of interest from viewing angles outside the input distribution. For instance, users may want to observe a scene from high-elevation angles, often missing from captured views. While novel view interpolation has seen significant advancements, this out-of-distribution novel view synthesis (OOD-NVS) task remains under-explored, in both evaluation protocols and methodology. A related research problem is 3D reconstruction from sparse input views, where methods often hallucinate unseen content (Liu et al., 2023a; Chan et al., 2023; Kwak et al., 2024; Liu et al., 2023b). While hallucination can be beneficial for creative applications, it may be undesirable in settings that demand accurate reconstructions, such as 3D visualization of surgical procedures (Hein et al., 2024), and unnecessary in typical daily capture scenarios.
@@
Figure 1: We investigate the out-of-distribution (OOD) novel view synthesis (NVS), where test views significantly differ from input views. This scenario contrasts with prior in-distribution NVS, where test views interpolate between densely captured input views, Sparse NVS with a few large-baseline input views, and Nerfbusters NVS (Warburg et al., 2023), where test views share similar angles with input views. Existing NVS methods, including MipNeRF360 (Barron et al., 2022), and those designed for sparse inputs like LaRa (Chen et al., 2024a), face challenges in this setting, while our method shows notable improvements.
@@
Unlike previous works that rely on hand-crafted regularization techniques (Xie et al., 2024a; Li et al., 2024b), we adapt point transformer (Zhao et al., 2021), an attention-based architecture designed for 3D scene understanding, to process the 3DGS as a point cloud set with Gaussian attributes serving as features. The attention mechanism in the point transformer learns to capture multi-view information embedded in the 3DGS, focusing on the local neighborhood within the spatial structure pre-computed by the initial 3DGS. It outputs residuals that are added to the input Gaussian attributes. The updated 3DGS is then rendered from novel views, and a photometric error between the rendered and ground-truth images is minimized to train the SplatFormer. We curate large-scale training pairs of initial, flawed 3DGS sets, and ground-truth images of in-distribution and OOD views using ShapeNet and Objaverse 1.0, which are made feasible by the fast optimization of 3DGS and the availability of large-scale 3D and multi-view datasets. By training on this dataset, SplatFormer learns generalizable priors for refining 3DGS, effectively removing artifacts in the OOD views while maintaining 3D consistency.
@@ 2 Related Work
Novel View Interpolation. In most NVS scenarios, both input and test views are sampled from the same distribution, typically following a fixed trajectory or a hemispherical pattern, as Blender NeRF (Mildenhall et al., 2020), LLFF (Mildenhall et al., 2019), and Phototourism (Jin et al., 2021). Seminal works like NeRF (Mildenhall et al., 2020), InstantNGP (Müller et al., 2022), and 3DGS (Kerbl et al., 2023) have demonstrated strong performance under this view interpolation protocol. However, as we will demonstrate later, they encounter difficulties in rendering novel views from out-of-distribution (OOD) angles, a challenge that remains less explored.
@@
Sparse View Reconstruction. Another line of research focuses on reconstructing 3D scenes from sparse set of input views, typically no more than four. Some approaches (Yu et al., 2021; Charatan et al., 2024; Wewer et al., 2024; Chen et al., 2024a; b; Fan et al., 2024b) directly predict 3D primitives from these sparse inputs. Others (Liu et al., 2023a; Sargent et al., 2024; Liu et al., 2023b; Kong et al., 2024; Yang et al., 2024; Abdal et al., 2024; Liu et al., 2024) finetune pretrained 2D diffusion models to generate novel views or repair degraded renderings. However, these methods are constrained by a limited number of input views and often rely on hallucinating unobserved regions. The sparse-view setting is well-suited for creative tasks but less applicable for accurate and faithful reconstructions when dense input views are available.
@@
Casually Captured Neural Radiance Fields. Unlike the standard interpolation setup, Nerfbusters (Warburg et al., 2023) captures input and test views along separate trajectories, closely relevant to the OOD-NVS problem we are addressing. However, their input and test views remain relatively similar in viewing angles, and the observed artifacts are primarily caused by the “invisibility issue”, where test views fall outside the input observation sphere, rather than from significant viewpoint deviations. In contrast, our approach tackles large viewpoint shifts without addressing invisibility, emphasizing generalization across substantial angle deviations.
@@
Regularization Techniques for Unconstrained Reconstruction. Sparse input views significantly degrade NVS performance, leading to various works exploring geometric priors, spatial regularity constraints, and data-driven priors. Geometric priors have been used in SuGaR (Guédon & Lepetit, 2024), 2DGS (Huang et al., 2024a), and GeoGaussian (Li et al., 2024b), which apply handcrafted self-supervision losses to better align Gaussian splats with surface geometry. Similarly, Spatial regularity constraints, explored in SplatFields (Mihajlovic et al., 2024) and ZeroRF (Shi et al., 2024), integrate Deep Image Prior (Ulyanov et al., 2018) to regularize 3DGS and NeRF reconstructions, producing more robust results from sparse inputs. However, these methods offer limited improvements as they do not leverage external data. Data-driven priors have been adopted in several works. FSGS (Zhu et al., 2024) and DNGaussian (Li et al., 2024a) supervise depth maps using deep stereo models but suffer from scale ambiguity. InstantSplat (Fan et al., 2024a) uses dense point clouds for 3DGS initialization, though it struggles with overfitting. Nerfbusters (Warburg et al., 2023) pretrains a diffusion model for post-processing NeRF, achieving only marginal improvements. Appearance priors (Zhu et al., 2024; Sargent et al., 2024; Wu et al., 2024a; Gao et al., 2024; Kwak et al., 2024) use 2D diffusion models to regularize novel view renderings, but often struggle with multi-view consistency. Additionally, SSDNeRF (Chen et al., 2023) and HypNeRF (Sen et al., 2023) pretrain 3D feature grids on object-centric datasets, yet underperform in multi-category experiments.
@@
Learning-based 2D-to-3D Models. Another special case of data-driven models involves training feed-forward models on large-scale multi-view image datasets to predict 3D representations from 2D images. Several methods (Liu et al., 2023b; Höllein et al., 2024; Kong et al., 2024; Liu et al., 2023a), et al. fine-tune pretrained diffusion models to generate multi-view images from one or a few input views. PixelNeRF (Yu et al., 2021), MVSplat (Chen et al., 2024b), and related works (Charatan et al., 2024; Wewer et al., 2024) transform 2D image features into NeRF or Gaussian splats. Although these models can learn generic priors from multi-view datasets, they are typically constrained to only a few input views, limiting their ability to fully leverage larger multi-view inputs.
@@
3D Point Processing Techniques are central to our work and widely used across 3D tasks. Unlike 2D image features or 3D grids, point clouds are unordered and unevenly distributed, requiring specialized architectures to handle their irregularity and sparsity. Solutions include sparse convolution (Choy et al., 2019), MLPs (Qi et al., 2017), and transformers (Wang, 2023; Yang et al., 2023; Zhao et al., 2021). The point transformer (Zhao et al., 2021; Wu et al., 2022; 2024b), using attention to model spatial relationships, has proven particularly effective. In 3D reconstruction, CVT-xRF (Zhong et al., 2024) employs a 3D transformer to predict ray point attributes based on local neighborhoods, while LSM (Fan et al., 2024b) utilizes a point transformer to predict 3DGS from two input views. Our work is the first to adapt the point transformer for refining 3D Gaussian splats (3DGS), leveraging its ability to capture spatial relationships in irregular point clouds to enhance novel view synthesis fidelity.
@@ 4 Robust Out-of-distribution Novel View Synthesis
Key Observation. As shown in Fig. 2, the reconstruction quality degrades significantly as the test camera’s elevation increases, highlighting a key limitation of 3DGS in handling OOD views. The challenge is to make the representation robust to such viewpoint changes while preserving the advantages of 3DGS, such as real-time rendering and compatibility with rasterization-based tools. Addressing this limitation by incorporating priors and constraints into the optimization of 3DGS is a complex task. Previous approaches have attempted to address this using geometric constraints (Huang et al., 2024a; Mihajlovic et al., 2024) and data-driven priors (Fan et al., 2024a). However, as demonstrated later (Tab. 1), these methods fall short in achieving robust novel-view synthesis, emphasizing the need for a more effective solution. We believe that solving this issue requires incorporating three key aspects: leveraging generic priors from large-scale datasets, ensuring 3D consistency in renderings, and fully utilizing rich geometric information from all input views.
@@ 4.1 Learning Data-Driven Prior
Dataset. To enable SplatFormer to refine imperfect Gaussian splats using a data-driven prior, we curated a large dataset containing pairs of Gaussian primitives and corresponding multi-view images. We utilized 33k and 48k scenes from the ShapeNet (Chang et al., 2015) and Objaverse-1.0 (Deitke et al., 2023) datasets respectively. These assets were rendered from low-elevation input views and high-elevation out-of-distribution (OOD) views. The initial splats were generated from the low-elevation views (following Sec. 3). The data collection process, which required approximately 3000 GPU hours, was efficiently executed using budget GPUs like the RTX-2080Ti. We released the data and corresponding rendering code to facilitate future research.
@@
Figure 4: Novel View Synthesis under Out-of-Distribution Camera Angles. The first column shows 4 out of 32 input views. Here, we compare our method with LaRa (Chen et al., 2024a), SplatFields (Mihajlovic et al., 2024), MipNeRF360 (Barron et al., 2022), 2DGS (Huang et al., 2024a), and 3DGS (Kerbl et al., 2023). Results on Objaverse-OOD evaluation scenes; a comprehensive comparison with all the baselines is provided in the appendix (Fig. H.4).
@@ 5 Experiments
We begin by outlining our experimental setup, followed by a description of the evaluation protocol and the baseline methods used for comparison. Next, we present the results on OOD-NVS, cross-dataset generalization, and ablation studies. Finally, we discuss the limitations of our approach and potential directions for future research.
@@
OOD-NVS. As visualized in Fig 1, for a centered object, we simulate an input camera capturing 360-degree azimuths at low elevations. The camera takes $N_{\textrm{in}}$ photos from evenly spaced azimuths, with its elevation following a sinusoidal pattern defined by frequency $f$ and maximal elevation $\phi_{\textrm{max}}$. This mimics a user recording the target with physical constraints preventing stable top-down captures. For OOD test views, we set their elevations $\phi_{\textrm{ood}}\gg\phi_{\textrm{max}}$, simulating a top-down perspective.
@@
Evaluation Sets. We use Blender to render 20 objects from ShapeNet (Chang et al., 2015) and Objaverse-v1 (Deitke et al., 2023) each. We select common objects and scenes with meaningful top-down views, such as city streets and buildings, avoiding those with large cavities that are invisible from low-elevation views. Due to varying object heights and shapes, we render two input camera trajectories with $\phi_{\textrm{max}}=10^{\circ}$ and $20^{\circ}$, creating two input-target splits to represent different levels of view deviation. So each scene has two experiments with different input views and the same target views. We average evaluation metrics across the two sets of experiments. Each input trajectory consists of $N_{\textrm{in}}=32$ views. The OOD test set includes $N_{\textrm{out}}=9$ views, uniformly distributed from the top sphere with $\phi_{\textrm{ood}}\geq 70^{\circ}$. All renderings are at a resolution of $256\times 256$.
@@
Baselines. We evaluate SplatFormer using the OOD-NVS protocol, comparing it with several publicly available state-of-the-art methods from each method category. First, we include per-scene approaches designed primarily for in-distribution NVS, including InstantNGP (Müller et al., 2022), 3DGS (Kerbl et al., 2023), and MipNeRF360 (Barron et al., 2022). Next, we examine regularized 3DGS variants without external priors, including 2DGS (Huang et al., 2024a) and SplatFields (Mihajlovic et al., 2024). We further consider methods which regularize per-scene optimization using external priors, including InstantSplat (Fan et al., 2024a), FSGS (Zhu et al., 2024), SSDNeRF (Chen et al., 2023), and Nerfbusters (Warburg et al., 2023). Finally, we examine feed-forward models, which directly produce 2D Gaussians (LaRa (Chen et al., 2024a)) or multi-view images (SyncDreamer  (Liu et al., 2023b) and EscherNet (Kong et al., 2024)) from input images.
@@
For a fair comparison, we retrained SSDNeRF, LaRa, SyncDreamer, and EscherNet using the same training sets as SplatFormer. For SyncDreamer, which only supports single-view input, we selected the highest-elevation view. For EscherNet, we use three randomly sampled input views during training and all input views during inference. For LaRa, which is limited to four input views due to memory constraints, we chose four large-baseline views to maximize scene coverage. More details are provided in appendix (Sec. G).
@@
Results on OOD-NVS. Qualitative results (Fig. 4) show that LaRa produces blurry outputs, while MipNeRF360 suffers from floater artifacts and SplatFields smooths out fine details. Both 2DGS and 3DGS exhibit spiky artifacts. In contrast, SplatFormer significantly reduces the artifacts present in 3DGS, completes surface reconstruction, and even restores certain geometric properties, such as interlaced structures. While our method still faces challenges with high-frequency texture details, it outperforms previous approaches in terms of fidelity and consistency in out-of-distribution views, which is also supported by the clear quantitative improvements demonstrated in our results (Tab. 1). Additional visual results are provided in the appendix (Sec. F) and the supplementary video.
@@
Figure 5: Cross-dataset Generalization. SplatFormer trained on Objaverse successfully mitigates artifacts in OOD views in the GSO (Downs et al., 2022) dataset and our real-world object-centric captures. Additional results are presented in the appendix (Fig. H.5 and Fig. H.6).
@@
On the GSO-OOD evaluation set, SplatFormer achieves substantial improvements in both metrics and visual quality. Even on the real-world dataset, despite being trained exclusively on synthetic data, SplatFormer reduces artifacts. These improvements are reflected in the SSIM and LPIPS metrics, though we observed rather minimal improvements in PSNR, which we attribute to the pixel-wise PSNR’s sensitivity to imperfect calibration and our method’s limitation in modeling specular effects. Our method also outperforms MipNeRF360 and 2DGS, the best-performing baselines in Objaverse-OOD (Tab. 1). Additionally, we evaluate Nerfbusters (Warburg et al., 2023), which addresses robust novel view synthesis from novel camera trajectories—a relevant challenge to OOD-NVS. However, we find that Nerfbusters tends to mistakenly remove key scene content as floaters, leading to incomplete geometry.
@@
DiffBIR-stage1 (Lin et al., 2021) 24.81 0.892 0.163
DiffBIR-stage2 (Lin et al., 2021) 24.24 0.858 0.174
@@
3D vs 2D Denoising. An alternative strategy for refining OOD-NVS renderings is to use 2D image restoration methods. To explore this, we use DiffBIR (Lin et al., 2021), a state-of-the-art image restoration method, to denoise 3DGS renderings. DiffBIR consists of two cascaded models: a first-stage image-to-image regressor to remove artifacts and a second-stage diffusion-based generator (Rombach et al., 2021) to in-paint missing details. We trained both stages using pairs of OOD 3DGS renderings and ground-truth images from our ShapeNet-OOD training set. To address multi-view inconsistencies in the denoised images, we also used the generated images to retrain the 3DGS. This experiment is similar to Sp2360 (Paul et al., 2024), which uses cascaded 2D diffusion priors to regularize 3DGS from sparse-view inputs. As shown in Tab. 5, while 2D denoising methods improve the original 3DGS, they significantly underperform compared to SplatFormer and fail to recover geometric details. See the appendix (Fig. C.1) for visual comparisons.